# One-loop corrections to the neutrino in the N-B-LSSM

### v1

@@ V Conclusion
In this paper, we systematically study the one-loop correction to the neutrino mass matrix in the N-B-LSSM and its fitting to experimental data on neutrino oscillations. We analyse the contributions to the neutrino mass from the loop diagrams of scalar lepton-chargino, scalar neutrino-neutralino and so on. Then we obtain the sum of tree-level and one-loop contributions to the neutrino mixing matrix. The one-loop corrected light neutrino effective mass matrix $M_{\nu}^{eff}$ is derived, and using the top-down approach, we give formulas for the neutrino mass variances and mixing angles. For the neutrino mass spectra, NO and IO conditions are discussed numerically. The light neutrino masses are very tiny and there are five experiment constraints (three mixing angles and two mass squared differences), the obtained suitable parameter space is narrow. The numerical results show that the model can fit the current experimental data, including the neutrino mass squared deviation and the mixing angle, well within a specific parameter range.
@@ Appendix A Mass matrix and coupling in N-B-LSSM
$m^{2}_{\tilde{e}}=\left(\begin{array}[]{cc}m_{\tilde{e}_{L}\tilde{e}_{L}^{*}}&\frac{1}{\sqrt{2}}v_{d}T_{e}^{\dagger}-\frac{1}{2}v_{u}{\lambda}v_{S}Y_{e}^{\dagger}\\ \frac{1}{\sqrt{2}}\sqrt{2}v_{d}T_{e}-\frac{1}{2}v_{u}{\lambda^{*}}v_{S}Y_{e}&m_{\tilde{e}_{R}\tilde{e}_{R}^{*}}\end{array}\right),$ (75)
@@
$Z^{E}m^{2}_{\tilde{e}}Z^{E,\dagger}=m^{diag}_{2,\tilde{e}}$ (77)
@@
$m^{2}_{\nu^{I}}=\left(\begin{array}[]{cc}m_{\sigma_{L}\sigma_{L}}&m_{\sigma_{R}\sigma_{L}}\\ m_{\sigma_{L}\sigma_{R}}&m_{\sigma_{R}\sigma_{R}}\end{array}\right)$ (78)
@@
$m^{2}_{\nu^{R}}=\left(\begin{array}[]{cc}m_{\phi_{L}\phi_{L}}&m_{\phi_{R}\phi_{L}}^{T}\\ m_{\phi_{L}\phi_{R}}&m_{\phi_{R}\phi_{R}}\end{array}\right)$ (81)
@@ Appendix B The definition of $F_{1},F_{2},F_{3},F_{4}$
$F_{2}(m_{1},m_{2})=\frac{1}{64\pi^{2}}\Big{(}-\frac{(m_{1}^{2}-3m_{2}^{2})}{(2m_{1}^{2}-2m_{2}^{2})}+\frac{(m_{1}^{4}-2m_{1}^{2}m_{2}^{2})\log{m_{1}^{2}}}{(m_{1}^{2}-m_{2}^{2})^{2}}$
@@
$F_{3}(m_{1},m_{2})=\frac{1}{32(m_{1}^{2}-m_{2}^{2})^{3}\pi^{2}}\Big{(}m_{1}^{4}-m_{2}^{4}-2m_{1}^{2}m_{2}^{2}\log{m_{1}^{2}}+2m_{1}^{2}m_{2}^{2}\log{m_{2}^{2}}\Big{)},$
$F_{4}(m_{1},m_{2})=\frac{1}{96(m_{1}^{2}-m_{2}^{2})^{4}\pi^{2}}(m_{1}^{6}-6m_{1}^{4}m_{2}^{2}+3m_{1}^{2}m_{2}^{4}+2m_{2}^{6}$
$\hskip 71.13188pt+6m_{1}^{2}m_{2}^{4}\log{m_{1}^{2}}-6m_{1}^{2}m_{2}^{4}\log{m_{2}^{2}}).$ (84)
@@ Appendix C Specific forms of the three mixing angles
$\left(\begin{array}[]{c}(U_{\nu})_{11}\\ (U_{\nu})_{21}\\ (U_{\nu})_{31}\end{array}\right)\;=\frac{1}{\sqrt{|X_{1}|^{2}+|Y_{1}|^{2}+|Z_{1}|^{2}}}\left(\begin{array}[]{c}X_{1}\\ Y_{1}\\ Z_{1}\end{array}\right)\;.$ (91)
$\left(\begin{array}[]{c}(U_{\nu})_{12}\\ (U_{\nu})_{22}\\ (U_{\nu})_{32}\end{array}\right)\;=\frac{1}{\sqrt{|X_{2}|^{2}+|Y_{2}|^{2}+|Z_{2}|^{2}}}\left(\begin{array}[]{c}X_{2}\\ Y_{2}\\ Z_{2}\end{array}\right)\;.$ (98)
@@
$X_{3}=\mathcal{H}_{12}\mathcal{H}_{23}-\mathcal{H}_{13}(\mathcal{H}_{22}-m_{\nu 3}^{2}),~{}~{}~{}~{}~{}~{}~{}~{}~{}Y_{3}=\mathcal{H}_{12}\mathcal{H}_{13}-\mathcal{H}_{23}(\mathcal{H}_{11}-m_{\nu 3}^{2})$
@@ Acknowledgements.
This work is supported by National Natural Science Foundation of China (NNSFC)(No.12075074), Natural Science Foundation of Hebei Province(A2020201002, A2022201022, A2022201017, A2023201040), Natural Science Foundation of Hebei Education Department (QN2022173), Post-graduate’s Innovation Fund Project of Hebei University (HBU2024SS042).

### v2

@@ V Conclusion
In this paper, we systematically study the one-loop correction to the neutrino mass matrix in the N-B-LSSM and its fitting to experimental data on neutrino oscillations. We analyse the contributions to the neutrino mass from the loop diagrams of scalar lepton-chargino, scalar neutrino-neutralino and so on. Then we obtain the sum of tree-level and one-loop contributions to the neutrino mixing matrix. The one-loop corrected light neutrino effective mass matrix $M_{\nu}^{eff}$ is derived, and using the ’top-down’ approach, we give formulas for the neutrino mass variances and mixing angles. For the neutrino mass spectra, NO and IO conditions are discussed numerically. The light neutrino masses are very tiny and there are five experiment constraints (three mixing angles and two mass squared differences), the obtained suitable parameter space is narrow. The numerical results show that the model can fit the current experimental data, including the neutrino mass squared deviation and the mixing angle, well within a specific parameter range.
@@ Appendix A Mass matrix and coupling in N-B-LSSM
$m^{2}_{\tilde{e}}=\left(\begin{array}[]{cc}m_{\tilde{e}_{L}\tilde{e}_{L}^{*}}&\frac{1}{\sqrt{2}}v_{d}T_{e}^{\dagger}-\frac{1}{2}v_{u}{\lambda}v_{S}Y_{e}^{\dagger}\\ \frac{1}{\sqrt{2}}v_{d}T_{e}-\frac{1}{2}v_{u}{\lambda^{*}}v_{S}Y_{e}&m_{\tilde{e}_{R}\tilde{e}_{R}^{*}}\end{array}\right),$ (84)
@@
$Z^{E}m^{2}_{\tilde{e}}Z^{E,\dagger}=m^{diag}_{2,\tilde{e}}.$ (86)
@@
$m^{2}_{\nu^{I}}=\left(\begin{array}[]{cc}m_{\sigma_{L}\sigma_{L}}&m_{\sigma_{R}\sigma_{L}}\\ m_{\sigma_{L}\sigma_{R}}&m_{\sigma_{R}\sigma_{R}}\end{array}\right),$ (87)
@@
$m^{2}_{\nu^{R}}=\left(\begin{array}[]{cc}m_{\phi_{L}\phi_{L}}&m_{\phi_{R}\phi_{L}}^{T}\\ m_{\phi_{L}\phi_{R}}&m_{\phi_{R}\phi_{R}}\end{array}\right),$ (90)
@@ Appendix B The definition of $F_{1},F_{2},F_{3},F_{4}$
$F_{2}(m_{1},m_{2})=\frac{1}{64\pi^{2}}\Big{(}-\frac{m_{1}^{2}-3m_{2}^{2}}{2m_{1}^{2}-2m_{2}^{2}}+\frac{(m_{1}^{4}-2m_{1}^{2}m_{2}^{2})\log{m_{1}^{2}}}{(m_{1}^{2}-m_{2}^{2})^{2}}$
@@
$F_{3}(m_{1},m_{2})=\frac{1}{32(m_{1}^{2}-m_{2}^{2})^{3}\pi^{2}}\Big{(}m_{1}^{4}-m_{2}^{4}-2m_{1}^{2}m_{2}^{2}\log{\frac{m_{1}^{2}}{m_{2}^{2}}}\Big{)},$
$F_{4}(m_{1},m_{2})=\frac{1}{96(m_{1}^{2}-m_{2}^{2})^{4}\pi^{2}}(m_{1}^{6}-6m_{1}^{4}m_{2}^{2}+3m_{1}^{2}m_{2}^{4}+2m_{2}^{6}+6m_{1}^{2}m_{2}^{4}\log{\frac{m_{1}^{2}}{m_{2}^{2}}}).$ (93)
@@ Appendix C Specific forms of the three mixing angles
$\left(\begin{array}[]{c}(U_{\nu})_{11}\\ (U_{\nu})_{21}\\ (U_{\nu})_{31}\end{array}\right)\;=\frac{1}{\sqrt{|X_{1}|^{2}+|Y_{1}|^{2}+|Z_{1}|^{2}}}\left(\begin{array}[]{c}X_{1}\\ Y_{1}\\ Z_{1}\end{array}\right)\;,$ (100)
$\left(\begin{array}[]{c}(U_{\nu})_{12}\\ (U_{\nu})_{22}\\ (U_{\nu})_{32}\end{array}\right)\;=\frac{1}{\sqrt{|X_{2}|^{2}+|Y_{2}|^{2}+|Z_{2}|^{2}}}\left(\begin{array}[]{c}X_{2}\\ Y_{2}\\ Z_{2}\end{array}\right)\;,$ (107)
@@
$X_{3}=\mathcal{H}_{12}\mathcal{H}_{23}-\mathcal{H}_{13}(\mathcal{H}_{22}-m_{\nu 3}^{2}),~{}~{}~{}~{}~{}~{}~{}~{}~{}Y_{3}=\mathcal{H}_{12}\mathcal{H}_{13}-\mathcal{H}_{23}(\mathcal{H}_{11}-m_{\nu 3}^{2}),$
@@ Acknowledgements.
This work is supported by National Natural Science Foundation of China (NNSFC) (No.12075074), Natural Science Foundation of Hebei Province (A2023201040, A2022201022, A2022201017, A2023201041), Natural Science Foundation of Hebei Education Department (QN2022173), Post-graduate’s Innovation Fund Project of Hebei University (HBU2024SS042), the Project of the China Scholarship Council (CSC) No. 202408130113. X. Dong acknowledges support from Fundação para a Ciência e a Tecnologia (FCT, Portugal) through the projects CFTP FCT Unit UIDB/00777/2020 and UIDP/00777/2020.